OpenStudy (anonymous):

Explain how you would find the zeros for f(x)=x^3+4x^2-x-10 How many zeros does this function have and how do you know

OpenStudy (anonymous):

degree of 3 means at most there are 3 solutions (X^3) set as 0=x^3+4x^2-x-10 use synthetic division ans=-2, -1 +/-[\sqrt{6}\]